# Adaptive Audio-Visual Speech Recognition via Matryoshka-Based Multimodal LLMs

### v1

@@ 3.1 Implementation Details
Llama-MTSK Details. We use Whisper Small and Medium (Radford et al., 2023) as pre-trained audio encoder, whilst AV-HuBERT Large (Shi et al., 2022) for computing the video tokens. Their weights remain frozen throughout the training phase. The projectors consist of two linear layers with ReLU activation in between. As for the LLM, based on the task and dataset, we experiment with $3$ base pre-trained models of varying size from the Llama 3 family (Dubey et al., 2024): Llama 3.1-8B, Llama 3.2-3B, and Llama 3.2-1B. Each LoRA module used to fine-tune the query and key projection matrices of each LLM’s self-attention layer has a bottleneck dimension $r$ such that the original LLM’s hidden size is reduced of a factor $32$ for Llama 3.2-3B and 3.2-1B, and $64$ for Llama 3.1-8B (e.g., for Llama 3.2-1B, since the hidden size is $2048$, the rank is set to $2048/32=64$). The hyperparameter $s$ is set to $\frac{1}{8}$.
Table 1: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
@@
[…]
Llama [Uncaptioned image] SS 2.3 2.2 3.3 3.6
@@
[…]
Llama [Uncaptioned image] SS 3.4 4.7 4.8 6.4
@@
Audio-Visual Token Compression Rates. We choose the audio and video compression rates to train and evaluate Llama-MTSK carefully, based on the studied tasks. For ASR, we apply compression rates in the range of {$4$, $8$, $12$, $16$, $20$}. For VSR, since the task is more challenging, we can afford smaller rates: {$1$, $2$, $3$, $4$, $5$} (we also include the case in which no compression is applied). For AVSR, we apply audio rates in {$4$, $16$} and video rates in {$2$, $5$}, leading to $4$ audio-visual configurations. To compress the audio and video tokens, either we apply average pooling with kernel size and stride equal to the desired compression rate, or we stack consecutive frames along the hidden dimension according to the rate (we denote this as “stacking”).
Table 2: Comparison between Llama [Uncaptioned image] and multiple SOTA methods on the LRS2 and LRS3 benchmarks. The “Lab. Hrs.” column with values X/Y specifies how many labeled hours have been used in training for LRS2 (X) and LRS3 (Y).
@@
[…]
Llama [Uncaptioned image] SS (4,2) 223/433 2.4 0.9
@@ 3.2 AVSR Main Results
Figure 3: WER results for the average pooling (left) and stacking (right) compression methods for the ASR task. We use Whisper Small as audio encoder and Llama 3.2-1B as LLM.
@@
We report the results achieved by Llama-MTSK MS, SS, and MSS on the LRS2 and LRS3 datasets in Table 1. We replace “MTSK” with [Uncaptioned image] in the tables and in the following sections to simplify the notation. For both datasets, we use Whisper Small as audio encoder. For the LLM, we use Llama 3.2-1B for LRS3 and Llama 3.2-3B for LRS2. The smaller size of the LRS2 dataset necessitates the larger LLM to mitigate higher WERs. We apply audio compression rates of $4$ and $16$ and video compression rates of $2$ and $5$, resulting in $4$ different compression configurations. We compare these results with those achieved by training Llama-AVSR independently on the $4$ configurations, leading to $4$ models. During inference, Llama-AVSR employs a separate model trained for each audio-video compression rate. In contrast, our Llama [Uncaptioned image] uses a single pre-trained model, activating the projector and LoRA [Uncaptioned image] modules corresponding to the desired compression rate. On the LRS3 dataset, the three proposed Llama [Uncaptioned image] approaches achieve comparable or superior performance to Llama-AVSR, particularly for the SS and MSS configurations. These two methods use LoRA modules specialized for specific compression rates, which are activated during inference based on specific requirements. On the LRS2 dataset, Llama [Uncaptioned image] SS outperforms all other approaches across all compression rates.
@@
Llama [Uncaptioned image] vs SOTA Methods. In Table 2, we compare Llama [Uncaptioned image] with state-of-the-art (SOTA) methods on LRS2 and LRS3 for the AVSR task. We equip Llama [Uncaptioned image] with Whisper Medium and Llama 3.1-8B. We report results from $5$ recent SOTA AVSR methods: CM-seq2seq (Ma et al., 2021), Efficient Conformer (Burchi and Timofte, 2023), auto-avsr (Ma et al., 2023b), Whisper-Flamingo (Rouditchenko et al., 2024), and USR (Haliassos et al., 2024a). Notably, all these methods do not reduce the token sequence length, whereas Llama-AVSR and Llama [Uncaptioned image] reduce the number of tokens by a factor $4$ for audio and $2$ for video. For LRS3, Llama [Uncaptioned image] achieves SOTA results, with its SS variant surpassing Llama-AVSR, which is trained on those specific compression rates, and outperforming methods like auto-avsr and USR, which use more training hours. For LRS2, Llama [Uncaptioned image] SS and MSS perform comparably to Llama-AVSR, while MS achieves better results. Additionally, our methods perform as well as or better than CM-seq2seq and Efficient Conformer but slightly underperform other SOTA methods. However, Llama [Uncaptioned image] is trained only on the $223$ hours of LRS2, whereas all competing methods utilize at least $1982$ hours. We leave for future work the integration of additional training data to enable a fairer comparison. Finally, more AVSR experiments can be found in the Appendix.
3.3 Additional Results
@@
Table 3: Comparison between Llama [Uncaptioned image] MS and a training-free Llama-AVSR-based approach that reduces the number of tokens via average pooling at inference time for the ASR task on the LRS3 dataset.
@@
[…]
ASR Results. For the ASR task, we consider $5$ compression rates in the range {$4$, $8$, $12$, $16$, $20$}. In Figure 3, we report the results on the LRS3 dataset when using average pooling compression (left) and stacking compression (right). With the exception of rate = $20$, all the three Llama [Uncaptioned image] methods outperform separately-trained Llama-AVSR methods. The MSS configuration achieves the best WER performance across all the compression rates, even surpassing or equaling the performance of Llama-AVSR trained at the lowest compression rate of $20$.
@@
VSR Results. Figure 4 shows WER results for the VSR task, similar to the ASR results in Figure 3. The video rates are {$1$, $2$, $3$, $4$, $5$}, lower than the ASR rates due to the greater complexity of VSR. For both average pooling and stack compression, all three Llama [Uncaptioned image] approaches outperform Llama-AVSR, with increasing gains at higher rates. The MS and SS approaches using average pooling achieve WER reductions exceeding $10$ at the highest rates. We attribute this improvement at higher compression rates to the joint training of multi-scale tokens. The performance of the three LoRA [Uncaptioned image] approaches varies slightly depending on the compression method, suggesting that no single approach is superior across all configurations. However, all of them significantly outperform Llama-AVSR.
Table 4: Computational cost analysis of Llama [Uncaptioned image] MS using different compression rates and Llama 3.1-8B.
(A,V) Rates # Tokens TFLOPs
(1,1) 757 11.40
[…]
(4,5) 182 2.74
(16,2) 163 2.46
(16,5) 88 1.33
@@
Llama [Uncaptioned image] vs Avg Pooling at Inference Time. Llama [Uncaptioned image] trains a single model that supports multiple scales at inference time by applying different compression rates. We compare our method with a training-free approach that trains a single Llama-AVSR model without compression and then applies the desired compression rate at inference on-the-fly by average pooling the tokens. In Table 3, we study the ASR setting with audio compression rates in the range {$2$, $4$, $6$, $8$, $10$}. The performance of the average-pooling baseline is severely impacted by a decrease in the number of tokens, while Llama [Uncaptioned image] MS is much more robust. These results demonstrate that Llama [Uncaptioned image] MS can be effectively used with diverse computational resources. Notably, even with limited resources, a compression rate of $8$ incurs minimal performance loss.
Computation Cost Analysis. Table 4 presents the computational benefits of using Llama [Uncaptioned image] . Specifically, we evaluate MS LoRA [Uncaptioned image] with Llama 3.1-8B as LLM and detail the associated inference costs. Without compression, we assume the LLM processes $500$ audio tokens, $250$ video tokens (the resolution of the audio encoder is twice that of the video encoder), and $7$ tokens for the textual prompt, totaling $757$. As shown in the table, our proposed approach yields speedups, reducing TFLOPs by over $8$x when applying compression rates of $16$ and $5$ for audio and video, respectively, thus substantially improving efficiency.
4 Conclusion
@@
This work introduces Llama-MTSK, a versatile audio-visual MLLM capable of elastic inference across multiple tasks and computational resources. Llama-MTSK exploits the concept of matryoshka representation learning to adapt the pre-trained LLM through ad-hoc LoRA [Uncaptioned image] modules, achieving performance comparable to or better than models separately trained on each compression rate while significantly reducing computational costs.
5 Limitations
[…]
@@ Appendix A Appendix
Figure 5: Additional WER results using stacking compression for the ASR task with {$2$, $4$, $6$, $8$, $10$} rates. We use the same configuration as in Figure 3.
A.1 Additional Experiments for ASR
In this section, we report additional results for the ASR task when using compression rates in a different range , specifically {$2$, $4$, $6$, $8$, $10$}. Compared to Figure 3, the increment between two consecutive rates is halved. We argue that it is more useful to use more diverse rates for ASR since we do not observe much deterioration of the WER results when doubling the rate (in Figure 5, the baseline Llama-AVSR achieves similar results when compressing the tokens of a factor $2$, $4$, and $6$). Figure 5 shows that Llama [Uncaptioned image] MS and MSS achieves comparable or better performance than Llama-AVSR. As for the SS approach, it performs slightly worse than Llama-AVSR for the first compression rates, and we believe this is because having a specific LoRA module for multiple rates which do not show WER deterioration leads to overfitting as one global LoRA is sufficient. This argument also explains why for rates $8$ and $10$ the MS variant performs better than the other ones.
@@ A.1 Additional Experiments for ASR
A.2 AVSR Results with Stacking Compression
@@
A.3 Full AVSR Results with Whisper Medium and LLama 3.1-8B
In Table 2, we only included for Llama-AVSR and Llama [Uncaptioned image] the results with audio and video compression rates equal to $4$ and $2$, respectively. In Table 5, we also report additional configurations of audio-video compression rates. We use Whisper medium as audio encoder and Llama 3.1-8B as LLM. Once more, our proposed methods perform on par or even better than independently-trained Llama-AVSR models for each compression rates configurations. In particular, we highlight the sizeable gains brought by all the three LoRA [Uncaptioned image] approaches for LRS3 when we apply the highest compression rates configuration (16,5).
Table 5: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. We employ Whisper medium and Llama 3.1-8B. Llama-AVSR trains $4$ independent models tailored to each configuration of audio-video compression rates.
@@
[…]
Llama [Uncaptioned image] SS 0.9 1.0 1.7 1.8
@@
[…]
Llama [Uncaptioned image] SS 2.4 2.1 2.9 2.9

### v2

@@ III-A Implementation Details
Llama-MTSK Details. We use Whisper Small and Medium [64] as pre-trained audio encoder, whilst AV-HuBERT Large [9] for computing the video tokens. Their weights remain frozen throughout the training phase. The projectors consist of two linear layers with ReLU activation in between. As for the LLM, based on the task and dataset, we experiment with $3$ base pre-trained models of varying size from the Llama 3 family [33]: Llama 3.1-8B, Llama 3.2-3B, and Llama 3.2-1B. Each LoRA module used to fine-tune the query and key projection matrices of each LLM’s self-attention layer has a bottleneck dimension $r$ such that the original LLM’s hidden size is reduced of a factor $32$ for Llama 3.2-3B and 3.2-1B, and $64$ for Llama 3.1-8B (e.g., for Llama 3.2-1B, since the hidden size is $2048$, the rank is $2048/32=64$). The hyperparameter $s$ is set to $\frac{1}{8}$.
TABLE I: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. Llama-AVSR trains 44 independent models tailored to each configuration of audio-video compression rates.
@@
Llama [Uncaptioned image] MS 2.6 2.7 3.7 4.1
Llama [Uncaptioned image] $S$ 2.3 2.2 3.3 3.6
@@
Llama [Uncaptioned image] MS 4.8 5.9 6.4 8.9
Llama [Uncaptioned image] $S$ 3.4 4.7 4.8 6.4
@@
Audio-Visual Token Compression Rates. We choose the audio and video compression rates to train and evaluate Llama-MTSK carefully, based on the studied tasks. For ASR, we apply compression rates in the range of {$4$, 88, $12$, $16$, 2020}. For VSR, since the task is more challenging, we can afford smaller rates: {$1$, 22, 33, 44, 55} (we also include the case in which no compression is applied). For AVSR, we apply audio rates in {$4$, 1616} and video rates in {$2$, 55}, leading to $4$ audio-visual configurations. To compress the audio and video tokens, either we apply average pooling with kernel size and stride equal to the desired compression rate, or we stack consecutive frames along the hidden dimension according to the rate (we denote this as “stacking”).
TABLE II: Comparison between Llama [Uncaptioned image] and multiple SOTA methods on the LRS2 and LRS3 benchmarks. The “Lab. Hrs.” column with values X/Y specifies how many labeled hours have been used in training for LRS2 (X) and LRS3 (Y).
@@
Llama-AVSR (4,2) 223/433 2.4 0.9
Llama [Uncaptioned image] MS (4,2) 223/433 2.1 1.0
Llama [Uncaptioned image] $S$ (4,2) 223/433 2.4 0.9
@@
Figure 3: WER results for the average pooling (left) and stacking (right) compression methods for the ASR task.
Training/Inference Details. Following [26, 63], we augment visual inputs through horizontal flipping, random cropping, and adaptive time masking, while for audio we only apply adaptive time masking. For training, we sample babble noise from the NOISEX dataset [65] using a uniform distribution. We define the textual prompts as in [26]: “Transcribe {task_prompt} to text.”, where task_prompt $\in$ {“speech”, “video”, “speech and video”}. We train our model for $10$ epochs with the AdamW optimizer with cosine annealing scheduler and weight decay set to $0.1$ using NVIDIA A40 GPUs. The learning rate is 1e-3 for ASR and AVSR tasks, and 5e-4 for VSR. For decoding, we use beam search with a beam width of $15$ and temperature of $0.6$. The evaluation metric for all the experiments is the Word Error Rate (WER, %).
III-B AVSR Main Results
@@
We report the results achieved by Llama-MTSK MS, SS, and MSS on the LRS2 and LRS3 datasets in Table I. We replace “MTSK” with [Uncaptioned image] in the tables and in the following sections to simplify the notation. For both datasets, we use Whisper Small as audio encoder. For the LLM, we use Llama 3.2-1B for LRS3 and Llama 3.2-3B for LRS2. The smaller size of the LRS2 dataset necessitates the larger LLM to mitigate higher WERs. We apply audio compression rates of $4$ and $16$ and video compression rates of $2$ and $5$, resulting in $4$ different compression configurations. We compare these results with those achieved by training Llama-AVSR independently on the $4$ configurations, leading to $4$ models. During inference, Llama-AVSR employs a separate model trained for each audio-video rate. In contrast, our Llama [Uncaptioned image] uses a single pre-trained model, activating the projector and LoRA [Uncaptioned image] modules corresponding to the desired rate. On the LRS3 dataset, the three proposed Llama [Uncaptioned image] approaches achieve comparable or superior performance to Llama-AVSR, particularly for the SS and MSS configurations. These two methods use LoRA modules specialized for specific rates, which are activated during inference based on specific requirements. On the LRS2 dataset, Llama [Uncaptioned image] SS outperforms all other approaches across all rates.
@@
Llama [Uncaptioned image] vs SOTA Methods. In Table II, we compare Llama [Uncaptioned image] with state-of-the-art (SOTA) methods on LRS2 and LRS3 for the AVSR task. We equip Llama [Uncaptioned image] with Whisper Medium and Llama 3.1-8B. We report results from $5$ recent SOTA AVSR methods: CM-seq2seq [5], Efficient Conformer [66], auto-avsr [63], Whisper-Flamingo [67], and USR [13]. Notably, all these methods do not reduce the token sequence length, whereas Llama-AVSR and Llama [Uncaptioned image] reduce the number of tokens by a factor $4$ for audio and $2$ for video. For LRS3, Llama [Uncaptioned image] achieves SOTA results, with its SS variant surpassing Llama-AVSR, which is trained on those specific compression rates, and outperforming methods like auto-avsr and USR, which use more training hours. For LRS2, Llama [Uncaptioned image] SS and MSS perform comparably to Llama-AVSR, while MS achieves better results. Additionally, our methods perform as well as or better than CM-seq2seq and Efficient Conformer but slightly underperform other SOTA methods. However, Llama [Uncaptioned image] is trained only on the $223$ hours of LRS2, whereas all competing methods utilize at least $1982$ hours. We leave for future work the integration of additional training data to enable a fairer comparison. Finally, more AVSR experiments can be found in the Appendix.
III-C Additional Results
@@
ASR Results. For the ASR task, we consider $5$ compression rates in the range {$4$, 88, $12$, $16$, 2020}. In Figure 3, we report the results on the LRS3 dataset when using average pooling compression (left) and stacking compression (right). With the exception of rate = $20$, all the three Llama [Uncaptioned image] methods outperform separately-trained Llama-AVSR methods. The MSS configuration achieves the best WER performance across all the compression rates, even surpassing or equaling the performance of Llama-AVSR trained at the lowest compression rate of $20$.
TABLE III: Comparison between Llama [Uncaptioned image] MS and a training-free Llama-AVSR-based approach that reduces the tokens via average pooling at inference time for the ASR task on LRS3.
@@
Llama [Uncaptioned image] MS 2.5 2.3 2.3 2.7 3.0
@@
VSR Results. Figure 4 shows WER results for the VSR task, similar to the ASR results in Figure 3. The video rates are {$1$, 22, 33, 44, 55}, lower than the ASR rates due to the greater complexity of VSR. For both average pooling and stack compression, all three Llama [Uncaptioned image] approaches outperform Llama-AVSR, with increasing gains at higher rates. The MS and SS approaches using average pooling achieve WER reductions exceeding $10$ at the highest rates. We attribute this improvement at higher compression rates to the joint training of multi-scale tokens. The performance of the three LoRA [Uncaptioned image] approaches varies slightly depending on the compression method, suggesting that no single approach is superior across all configurations. However, all of them significantly outperform Llama-AVSR.
@@
Llama [Uncaptioned image] vs Avg Pooling at Inference Time. Llama [Uncaptioned image] trains a single model that supports multiple scales at inference time by applying different compression rates. We compare our method with a training-free approach that trains a single Llama-AVSR model without compression and then applies the desired compression rate at inference on-the-fly by average pooling the tokens. In Table III, we study the ASR setting with audio rates in the range {$2$, 44, 66, 88, 1010}. The performance of the average-pooling baseline is severely impacted by a decrease in the number of tokens, while Llama [Uncaptioned image] MS is much more robust. These results demonstrate that Llama [Uncaptioned image] MS can be effectively used with diverse computational resources.
Computation Cost Analysis. In Figure 5, we illustrate the benefits of Llama [Uncaptioned image] MS in terms of TFLOPS and inference costs. Without compression (i.e., (1,1)), we assume the LLM processes $500$ audio tokens, $250$ video tokens (the resolution of the audio encoder is twice that of the video encoder), and $7$ tokens for the textual prompt, totaling $757$. By increasing the audio-visual compression rates, we reduce the number of tokens processed by the LLM, and thus the TFLOPs, by up to 8.68.6x when applying compression rates of (16,5), resulting in $88$ tokens. Despite this substantial reduction in TFLOPs, the resulting increase in WER remains modest. Moreover, Llama [Uncaptioned image] MS enables elastic inference by allowing users to select compression rates based on their computational constraints.
Figure 5: Comparison of Llama Refer to caption MS in terms of number of audio-visual processed tokens, WER, and TFLOPS.
IV Conclusion
@@
We introduce Llama-MTSK, a versatile audio-visual MLLM capable of elastic inference across multiple tasks and computational resources. Llama-MTSK exploits matryoshka representation learning to adapt the pre-trained LLM through ad-hoc LoRA [Uncaptioned image] modules, achieving performance comparable to or better than models separately trained on each compression rate while significantly reducing computational costs.
@@ Appendix A Appendix
Figure 6: Additional WER results using stacking compression for the ASR task with {22, 44, 66, 88, 1010} rates.
Appendix B Additional Experiments for ASR
@@
In this section, we report additional results for the ASR task when using compression rates in a different range, specifically {$2$, 44, 66, 88, 1010}. Compared to Figure 3 in the main paper, the increment between two consecutive rates is halved. We argue that it is more useful to use more diverse rates for ASR since we do not observe much deterioration of the WER results when doubling the rate (in Figure 6, the baseline Llama-AVSR achieves similar results when compressing the tokens of a factor $2$, 44, and $6$). Figure 6 shows that Llama [Uncaptioned image] MS and MSS achieves comparable or better performance than Llama-AVSR. As for the SS approach, it performs slightly worse than Llama-AVSR for the first compression rates, and we believe this is because having a specific LoRA module for multiple rates which do not show WER deterioration leads to overfitting as one global LoRA is sufficient. This argument also explains why for rates $8$ and $10$ the MS variant performs better than the other ones.
Appendix C AVSR Results with Stacking Compression
@@
Appendix D Full AVSR Results with Whisper Medium and LLama 3.1-8B
In Table 2 in the main paper, we only included for Llama-AVSR and Llama [Uncaptioned image] the results with audio and video compression rates equal to $4$ and $2$, respectively. In Table IV, we also report additional configurations of audio-video compression rates. We use Whisper medium as audio encoder and Llama 3.1-8B as LLM. Once more, our proposed methods perform on par or even better than independently-trained Llama-AVSR models for each compression rates configurations. In particular, we highlight the sizeable gains brought by all the three LoRA [Uncaptioned image] approaches for LRS3 when we apply the highest compression rates configuration (16,5).
TABLE IV: Comparison between Llama-AVSR and our proposed Llama [Uncaptioned image] MS, SS, and MSS approaches on LRS2 and LRS3 benchmarks. We employ Whisper medium and Llama 3.1-8B. Llama-AVSR trains 44 independent models tailored to each configuration of audio-video compression rates.
@@
Llama [Uncaptioned image] MS 1.0 1.1 1.5 1.6
Llama [Uncaptioned image] $S$ 0.9 1.0 1.7 1.8
@@
Llama [Uncaptioned image] MS 2.1 2.3 2.9 3.2
Llama [Uncaptioned image] $S$ 2.4 2.1 2.9 2.9